A production lot of size 100 is known to be 55% defective. A random sample of 12

items is selected without replacement. Determine the probability that there will be

no defective in the sample.


1
Expert's answer
2021-08-16T10:50:23-0400

Solution:

Required probability"=\dfrac{45}{100}\times \dfrac{44}{99}\times \dfrac{43}{98}\times ...\times\dfrac{34}{89}=0.00002738"